# Adaptive Cache Enhancement for Test-Time Adaptation of Vision-Language Models

### v1

@@ Abstract
Vision-language models (VLMs) exhibit remarkable zero-shot generalization but suffer performance degradation under distribution shifts in downstream tasks, particularly in the absence of labeled data. Test-Time Adaptation (TTA) addresses this challenge by enabling online optimization of VLMs during inference, eliminating the need for annotated data. Cache-based TTA methods exploit historical knowledge by maintaining a dynamic memory cache of low-entropy or high-confidence samples, promoting efficient adaptation to out-of-distribution data. Nevertheless, these methods face two critical challenges: (1) unreliable confidence metrics under significant distribution shifts, resulting in error accumulation within the cache and degraded adaptation performance; and (2) rigid decision boundaries that fail to accommodate substantial distributional variations, leading to suboptimal predictions. To overcome these limitations, we introduce the Adaptive Cache Enhancement (ACE) framework, which constructs a robust cache by selectively storing high-confidence or low-entropy image embeddings per class, guided by dynamic, class-specific thresholds initialized from zero-shot statistics and iteratively refined using an exponential moving average and exploration-augmented updates. This approach enables adaptive, class-wise decision boundaries, ensuring robust and accurate predictions across diverse visual distributions. Extensive experiments on 15 diverse benchmark datasets demonstrate that ACE achieves state-of-the-art performance, delivering superior robustness and generalization compared to existing TTA methods in challenging out-of-distribution scenarios.
Introduction
Vision-language models (VLMs) such as CLIP (Radford et al. 2021a) and ALIGN (Jia et al. 2021) exhibit remarkable zero-shot abilities in a wide range of tasks, including classification, retrieval, and segmentation (Zhou et al. 2021; Khattak et al. 2022; Xie et al. 2023; Iscen et al. 2024; Jiahaoli et al. 2024; Zhu et al. 2023). These models are trained to learn the semantic relationship between visual and textual information and can be used for image-text matching during inference. However, their effectiveness often decreases when handling unlabeled test data from domains that substantially differ from their training distribution (Agarwal et al. 2021; Menon, Chandratreya, and Vondrick 2024; Parashar et al. 2024).
Figure 1: Noisy CLIP predictions on different views of “dog” image. CLIP can be either overconfident on simple view but fail on strong augmentation.
@@ Introduction
Prompt-based TTA methods (Shu et al. 2022a; Feng et al. 2023a; Hassan et al. 2023) reduce the prediction entropy of VLMs by optimizing domain-specific prompts. However, their prompting optimization requires repeated backpropagation through the encoders of VLMs, resulting in substantial computational cost and hence limiting their scalability in resource-limited settings. Cache-based TTA strategies (Karmanov et al. 2024a; Zhang et al. 2024b, c, a) dynamically assemble a cache of high-confidence samples guided by VLM prediction entropy metrics, thus fine-tuning predictions with minimal parameter changes. For example, TDA (Karmanov et al. 2024a) uses adaptive caching to enhance the calibration of predictions, while DMN (Zhang et al. 2024c) utilizes short- and long-term memory banks to leverage past insights. Similarly, DPE (Zhang et al. 2024a) improves multimodal alignment through dual-modal residual learning.
However, as demonstrated in Figure 1, CLIP predictions vary significantly under different augmentations, making the entropy-based sample selection unreliable and potentially hindering the adaptability of the model (Khan and Fu 2024; Farina et al. 2024). Furthermore, the decision boundaries of the VLMs in target domains become inconsistent. Even with TTA, these models frequently fail to adjust the decision boundaries to accommodate domain shifts, resulting in fluctuating predictions. Although several methods (Zhang et al. 2024c, a) attempt to overcome this issue by using mean textual prototypes from enriched prompts, they are limited in handling the variability from intra-class diversity shifts.
To address the above challenges, we propose the Adaptive Cache Enhancement (ACE) framework, which constructs a robust cache by selectively storing high-confidence or low-entropy image embeddings for each class, directed by dynamic, class-specific thresholds initialized from zero-shot statistics and iteratively refined with an exponential moving average and exploration-augmented updates. By excluding over-confident samples from the cache, our method avoids overfitting, allows for adaptable class-specific decision boundaries, and ensures reliable and precise predictions across a broad range of data distributions. To this end, we make the following contributions in our work.
@@
Related Work
Vision-Language Models (VLMs)
@@
Large-scale pre-trained VLMs such as CLIP (Radford et al. 2021a) have demonstrated substantial proficiency in acquiring transferable representations that span both imagery and text. Extensive research has been conducted to optimize these models for downstream applications. Initial efforts, including prompt learning strategies (Zhou et al. 2021, 2022a), achieve notable success by implementing minimal supervision in few-shot learning environments. Recent methods, e.g., Tip-Adapter (Zhang et al. 2021) and TaskRes (Yu et al. 2023), have effectively utilized feature memories derived from a limited number of labeled instances for rapid adaptation. These memory-based approaches have gained popularity due to their ability to facilitate quick domain adaptation with minimal computational burden, eliminating the requirement for full model fine-tuning. Nevertheless, their dependency on labeled target-domain data limits their applicability in real-world contexts, where such data are often not available. Our research focuses on test-time adaptation, which compels VLMs to acclimate to new domains during the inference stage without access to labeled data. This scenario closely reflects the real-world setting, advancing the development of flexible vision-language systems.
Test-Time Adaptation (TTA) for VMLs
TTA enhances the adaptability of VLMs to unfamiliar target domains by exclusively leveraging unlabeled test data (Shu et al. 2022a; Ma et al. 2023; Farina et al. 2024; Zhao et al. 2024). Current methods are generally divided into prompt-based and cache-based categories. Prompt-based techniques refine textual embeddings throughout inference. For example, TPT (Shu et al. 2022a) refines prompts to maintain consistency across multiple augmented image views. DiffTPT (Feng et al. 2023a) applies diffusion-driven augmentation techniques to enhance view diversity, thus improving the adaptability of VLMs to unfamiliar test data. Conversely, cache-based methods utilize historical, instance-specific data extracted from test samples. Specifically, TDA (Karmanov et al. 2024a) creates caches containing positive and negative features to improve CLIP’s predictions through aggregated similarity measurements. DMN (Zhang et al. 2024c) combines static and dynamic memory networks, employing an advanced interaction mechanism to effectively leverage both few-shot and accumulated unlabeled data for robust adaptation. More recently, DPE (Zhang et al. 2024a) refines visual and textual prototypes in parallel to capture domain-specific semantics more precisely. Although prompt-based techniques are appreciated for their simplicity, cache-based strategies often achieve superior adaptability and computational efficiency by modeling target domain distributions through lightweight, memory-based mechanisms at test time. Expanding on these cache-based methods, our research specifically aims at improving cache quality through reliable sample selection and enhancing predictive strength by adaptively adjusting decision boundaries.
@@
Figure 2: Overview of the ACE Method. We introduce the Class-wise Adaptive Threshold Enhancement module, which refines the threshold online to allow more correct samples but low-confidence samples to be cached. By simultaneously minimizing entropy loss for the prototype residuals, and creating highly reliable caches, ACE enhances the overall multimodal generalization and robustness.
Proposed Method
@@
We propose an effective and reliable TTA strategy that includes a cache data structure, prototype residual learning modules, a flexible class-wise thresholding mechanism, and an adaptive threshold refinement module based on the statistic of pseudo-labels. We realize our method with CLIP (Radford et al. 2021a), a state-of-the-art VLM model, and DPE (Zhang et al. 2024a), a prototype residual vector learning technique.
Preliminaries
Zero-shot CLIP Prediction.
@@
Cache-based TTA.
Inspired by the key-value storage framework introduced in TIP-Adapter (Zhang et al. 2022), several cache-based TTA approaches (Karmanov et al. 2024a; Zhang et al. 2024a, b) have been proposed, which dynamically store and update high-confidence feature-label pairs during inference. The cache is organized as a priority queue, where the priority of each entry is determined by the entropy of CLIP’s predictions.
Let $\mathbf{W}_{\mathrm{CLIP}}=\left[\textbf{t}^{1},\textbf{t}^{2},\ldots,\textbf{t}^{C}\right]^{\top}$ be a stack of text embeddings for all classes. Each cache item is represented as:
@@
As test samples arrive sequentially, new entries are inserted into cache slots corresponding to their pseudo-labels. When the cache exceeds its capacity, entries with the highest entropy are evicted. During inference, a test feature $\mathbf{z}$ retrieves related cache items, and the final prediction is calculated by combining the CLIP logits with those derived from the cache:
@@
Following the DPE approach Zhang et al. (2024a), the feature for each cache slot, $\mathbf{F}_{\text{cache}}$, is computed as the mean of all stored features for that class, forming increasingly compact and representative class-specific prototypes as more data accumulates. The cache item is then constructed as:
@@
Adaptive Cache Enhancement (ACE)
Cache-based strategies demonstrate notable effectiveness in various domains. However, they are grappling with two substantial issues. First, entropy is not a reliable indicator when the data distribution changes. Second, fixed decision boundaries are not adaptive to distributional changes. To address these challenges, we introduce two specialized modules that specifically address each issue while synergistically strengthening each other as illustrated in Figure 2.
Curriculum Thresholding
The task of dynamically setting class-specific thresholds in alignment with the evolving learning progress of the model is challenging. Ideally, the threshold for each class can be set specifically based on evaluation metrics, e.g., recognition accuracy or entropy as,
$\mathcal{T}_{t}(c)=m_{t}(c)\cdot\mathcal{T}_{0}(c),$ (6)
where $\mathcal{T}_{t}(c)$ is the adaptive threshold for class $c$ at time step $t$, and $m_{t}(c)$ denotes the corresponding metric, whether it is accuracy or entropy. This approach ensures that classes with poorer performance, indicated by lower accuracy rates, are designated smaller thresholds (or conversely, higher entropy leads to larger thresholds), thereby permitting a greater number of their instances to be retained in the cache during the testing phase. However, dynamically adjusting these thresholds throughout the training process requires continuous accuracy estimates at time increments $t$, which impedes training efficiency.
In this study, we introduce Curriculum Thresholding, a thresholding mechanism that evaluates the learning status of classes without introducing any additional inference steps and/or requiring extra annotations. Central to our hypothesis is the postulation that when subjected to a high threshold, the quantity of confidently classified predictions ascribed to a particular class serves as a robust indicator of the learning progress for that class. Specifically, we proceed with the following steps:
$i\noindent\sigma_{t}(c)$ $=\sum_{n=1}^{N}\mathds{1}\left(\max\left(p(y|x)\right)\geq\mathcal{T}_{t-1}(c)\right)\cdot\mathds{1}\left(\arg\max(p(y|x))=c\right),$
or $=\sum_{n=1}^{N}\mathds{1}\left(\min\left(e(y|x)\right)\leq\mathcal{T}_{t-1}(c)\right)\cdot\mathds{1}\left(\arg\max\left(p(y|x)\right)=c\right),$
where $\sigma_{t}(c)$ measures the number of samples confidently predicted as class $c$ at time $t$, and $p(y|x)$ and $e(y|x)$ are the probability and entropy of the model to sample $x$.
We then normalize $\sigma_{t}(c)$ as,
$\sigma_{t}(c)\leftarrow\frac{\sigma_{t}(c)}{\max_{c}\sigma_{t}}$ (7)
@@
Online Thresholding Adaptation
@@
CLIP-ResNet-50 (Radford et al. 2021b) 58.16 21.83 51.41 56.15 33.37 44.18 40.69
@@
CoOp (Zhou et al. 2022b) 63.33 23.06 55.40 56.60 34.67 46.61 42.43
TPT (Shu et al. 2022b) 60.74 26.67 54.70 59.11 35.09 47.26 43.89
DiffTPT (Feng et al. 2023b) 60.80 31.06 55.80 58.80 37.10 48.71 45.69
TDA (Karmanov et al. 2024b) 61.35 30.29 55.54 62.58 38.12 49.58 46.63
TPS (Sui, Wang, and Yeung-Levy 2024) 61.47 30.48 54.96 62.87 37.14 49.38 46.36
DMN-ZS (Zhang et al. 2024d) 63.87 28.57 56.12 61.44 39.84 49.97 46.49
DPE (Zhang et al. 2024a) 63.41 30.15 56.72 63.72 40.03 50.81 47.66
CRG (Zhai et al. 2025) 65.26 29.69 56.07 - - - -
ACE-Probability 63.75 30.09 56.84 63.95 40.15 50.94 47.76
ACE-Entropy 63.70 30.20 56.89 63.82 40.25 50.97 47.79
CLIP-ViT-B/16 (Radford et al. 2021b) 66.73 47.87 60.86 73.98 46.09 59.11 57.20
@@
CoOp (Zhou et al. 2022b) 71.51 49.71 64.20 75.21 47.99 61.72 59.28
TPT (Shu et al. 2022b) 68.98 54.77 63.45 77.06 47.94 62.44 60.81
DiffTPT (Feng et al. 2023b) 70.30 55.68 65.10 75.00 46.80 62.28 60.52
TDA (Karmanov et al. 2024b) 69.51 60.11 64.67 80.24 50.54 65.01 63.89
TPS (Sui, Wang, and Yeung-Levy 2024) 70.19 60.08 64.73 80.27 49.95 65.04 63.76
DMN-ZS (Zhang et al. 2024d) 72.25 58.28 65.17 78.55 53.20 65.49 63.80
DPE (Zhang et al. 2024a) 71.91 59.63 65.44 80.40 52.26 65.93 64.43
CRG (Zhai et al. 2025) 75.01 63.67 64.66 - - - -
ACE-Probability 71.93 59.71 65.46 80.68 52.60 66.08 64.61
ACE-Entropy 71.93 59.62 65.51 80.72 52.66 66.09 64.63
@@
Test-Time Logit Inference and Loss Function.
We adopt DPE, the residual learning method proposed by Zhang et al. (2024d), to define textual and visual embedding prototypes as $\mathbf{t}_{c}=\frac{1}{S}\sum_{i=1}^{S}\mathcal{E}_{t}(\mathcal{P}^{c}_{i})$ and $\mathbf{v}_{c}=\frac{1}{S_{c}}\sum_{m=1}^{S_{c}}\mathbf{z}^{c}_{m}$. Here, $S$ is the number of prompts, $S_{c}\leq M$ is the total count of image features stored in the queue corresponding to class $c$, and $\mathbf{z}$ is the image features of $x$ that surpass the threshold $\mathcal{T}$. Once the two sets of multi-modal prototypes have been adjusted based on the last test sample, they are used to serve as the initial configuration for further refinement with the current test input. The preliminary outcomes derived from these steps are subsequently utilized to fine-tune the prototypes, making them adapted to $x$ as,
@@
where $\hat{\mathbf{t}}_{c}\in\mathbb{R}^{C\times d}$ and $\hat{\mathbf{v}}_{c}\in\mathbb{R}^{C\times d}$ are the learnable residual parameters that are initialized as zero and are used to optimize the prototypes for each given test input $x$. Similarly to Zhang et al. (2024d), we optimize these residual parameters to promote consistent predictions in a total of $N$ different augmented views of each given test image $x$ using the objective of unsupervised entropy minimization,
@@
where $\mathcal{A}_{n}(x)$ is an augmented view of $x$.
We adopt the alignment loss in DPE,
@@
CoOp (Zhou et al. 2021) 15.12 86.53 55.32 37.29 26.20 61.55 75.59 87.00 58.15 59.05 56.18
TPT (Shu et al. 2022a) 17.58 87.02 58.46 40.84 28.33 62.69 74.88 84.49 61.46 60.82 57.66
DiffTPT (Feng et al. 2023c) 17.60 86.89 60.71 40.72 41.04 63.53 79.21 83.40 62.72 62.67 59.85
TDA (Karmanov et al. 2024a) 17.61 89.70 57.78 43.74 42.11 68.74 77.75 86.18 62.53 64.18 61.03
DPE (Zhang et al. 2024a) 19.80 90.83 59.26 50.18 41.67 67.60 77.83 85.97 64.23 61.98 61.93
CRG (Zhai et al. 2025) 18.09 90.12 57.92 51.89 46.80 71.09 75.76 85.91 63.11 63.58 62.42
@@
CoOp (Zhou et al. 2021) 18.47 93.70 64.51 41.92 46.39 68.71 85.30 89.14 64.15 66.55 63.88
TPT (Shu et al. 2022a) 24.78 94.16 66.87 47.75 42.44 68.98 84.67 87.79 65.50 68.04 65.10
DiffTPT (Feng et al. 2023c) 25.60 92.49 67.01 47.00 43.13 70.10 87.23 88.22 65.74 62.67 65.47
TDA (Karmanov et al. 2024a) 23.91 94.24 67.28 47.40 58.00 71.42 86.14 88.63 67.62 70.66 67.53
DPE (Zhang et al. 2024a) 28.95 94.81 67.31 54.20 55.79 75.07 86.17 91.14 70.07 70.44 69.40
CRG (Zhai et al. 2025) 26.58 93.57 66.89 56.38 59.81 75.94 85.95 91.20 68.36 70.31 69.50

### v2

@@ Abstract
Vision-language models (VLMs) demonstrate strong zero-shot generalization but degrade under distribution shifts in unlabeled downstream tasks. Test-Time Adaptation (TTA) enables online optimization during inference without annotations. Cache-based TTA methods maintain dynamic caches of low-entropy or high-confidence samples for efficient out-of-distribution adaptation. However, they encounter two key issues: (1) unreliable confidence metrics causing cache errors and reduced performance; and (2) inflexible decision boundaries failing to handle distributional variations. To address these, we propose the Adaptive Cache Enhancement (ACE) framework, which builds a robust cache by selectively storing high-confidence or low-entropy embeddings per class. It employs dynamic, class-specific thresholds initialized from zero-shot statistics and refined via exponential moving average and exploration-augmented updates, enabling adaptive class-wise boundaries for accurate predictions across diverse distributions. Experiments on 15 benchmarks show ACE achieves state-of-the-art results, outperforming existing TTA methods in challenging out-of-distribution scenarios.
1 Introduction
Vision-language models (VLMs) such as CLIP [CLIP] and ALIGN [jia2021scaling] exhibit remarkable zero-shot abilities in a wide range of tasks, including classification, retrieval, and segmentation [CoOp, MaPLe, raclip_retrieval, reco_retrieval, rpn_seg, seg2]. These models are trained to learn the semantic relationship between visual and textual information and can be used for image-text matching during inference. However, their effectiveness often decreases when handling unlabeled test data from domains that substantially differ from their training distribution [vlms1, vlms2, vlms3].
Figure 1: Noisy CLIP predictions on different views of a “dog“. CLIP can be overconfident on the simple views and fail on others.
@@ 1 Introduction
Prompt-based TTA methods [tpt, difftpt, promptalign] reduce the prediction entropy of VLMs by optimizing domain-specific prompts. However, their prompting optimization requires repeated backpropagation through the encoders of VLMs, resulting in substantial computational cost and hence limiting their scalability in resource-limited settings. Cache-based TTA strategies [tda, boostadapter, dmn, dpe] dynamically assemble a cache of high-confidence samples guided by VLM prediction entropy metrics, thus fine-tuning predictions with minimal parameter changes. For example, TDA [tda] uses adaptive caching to enhance the calibration of predictions, while DMN [dmn] utilizes short- and long-term memory banks to leverage past insights. Similarly, DPE [dpe] improves multimodal alignment through dual-modal residual learning.
However, as demonstrated in Figure 1, CLIP predictions vary significantly under different augmentations, making the entropy-based sample selection unreliable and potentially hindering the adaptability of the model [entropy_unreliable, zero]. Furthermore, the decision boundaries of the VLMs in target domains become inconsistent. Even with TTA, these models frequently fail to adjust the decision boundaries to accommodate domain shifts, resulting in fluctuating predictions. Although several methods [dmn, dpe] attempt to overcome this issue by using mean textual prototypes from enriched prompts, they are limited in handling the variability from intra-class diversity shifts.
To address the above challenges, we propose the Adaptive Cache Enhancement (ACE) framework, which constructs a robust cache by selectively storing high-confidence or low-entropy image embeddings for each class, directed by dynamic, class-specific thresholds initialized from zero-shot statistics and iteratively refined with an exponential moving average and exploration-augmented updates. By excluding overconfident samples from the cache, our method avoids overfitting, allows for adaptable class-specific decision boundaries, and ensures reliable and precise predictions across a broad range of data distributions. To this end, we make the following contributions in our work.
@@
2 Related Work
2.1 Vision-Language Models (VLMs)
@@
Large-scale pre-trained VLMs such as CLIP [CLIP] have demonstrated substantial proficiency in acquiring transferable representations that span both imagery and text. Extensive research has been conducted to optimize these models for downstream applications. Initial efforts, including prompt learning strategies [CoOp, cocoop], achieve notable success by implementing minimal supervision in few-shot learning environments. Recent methods, e.g., Tip-Adapter [tipadapter] and TaskRes [taskres], have effectively utilized feature memories derived from a limited number of labeled instances for rapid adaptation. These memory-based approaches have gained popularity due to their ability to facilitate quick domain adaptation with minimal computational burden, eliminating the requirement for full model fine-tuning. Nevertheless, their dependency on labeled target-domain data limits their applicability in real-world contexts, where such data are often not available. Our research focuses on test-time adaptation, which compels VLMs to acclimate to new domains during the inference stage without access to labeled data. This scenario closely reflects the real-world setting, advancing the development of flexible vision-language systems.
2.2 Test-Time Adaptation (TTA) for VMLs
Test-time adaptation (TTA) improves vision-language models (VLMs) adaptability to unfamiliar domains using only unlabeled test data [tpt, ma2024swapprompt, zero, rlcf]. Methods fall into prompt-based and cache-based categories. Prompt-based approaches refine textual embeddings during inference, such as TPT [tpt] for augmented view consistency and DiffTPT [difftpt] for diffusion-augmented diversity. In contrast, cache-based methods leverage historical test samples, including TDA [tda] for positive/negative feature caches, DMN [dmn] for static/dynamic memory integration, and DPE [dpe] for parallel prototype refinement. While prompt-based methods are simple, cache-based ones offer better adaptability and efficiency via lightweight domain modeling. Recent advances include BCA [zhou2025bayesian] (Bayesian priors with overhead), FreeTTA [dai2025free] (training-free online EM/GMM), COSMIC [huang2025cosmic] (semantic cliques), and AWT [zhu2024awt] (augmentation-weighting-transport), yet they neglect class-specific threshold dynamics under shifts. Building on cache-based techniques, our work enhances cache quality through reliable selection and predictive strength via adaptive decision boundaries; ACE distinctively employs EMA-refined thresholding to boost reliability and intra-class diversity handling.
@@ 2.2 Test-Time Adaptation (TTA) for VMLs
Figure 2: Overview of the ACE Method. We introduce the Class-wise Adaptive Threshold Enhancement module, which refines the threshold online to allow more correct samples but low-confidence samples to be cached. By simultaneously minimizing entropy loss for the prototype residuals and creating highly reliable caches, ACE enhances the overall multimodal generalization and robustness.
Method Key Mechanism
Adaptive
Thresholds
Class
Specific
Efficiency
Focus
BCA Bayesian priors No Partial No
FreeTTA Online EM/GMM No No Yes*
COSMIC Semantic cliques No No Yes
AWT Augment/Weight/Transport No No Partial
ACE (Ours)
Curriculum
+ Exploration Thresholding
Yes Yes Yes
Table 1: Methodology comparison. ’*’ denotes training-free.
3 Proposed Method
@@
We propose an effective and reliable TTA strategy that includes a cache data structure, prototype residual learning modules, a flexible class-wise thresholding mechanism, and an adaptive threshold refinement module based on the statistic of pseudo-labels. We realize our method with CLIP [CLIP], a state-of-the-art VLM model, and DPE [dpe], a prototype residual vector learning technique.
3.1 Preliminaries
3.1.1 Zero-shot CLIP Prediction.
@@ 3.1 Preliminaries
3.1.2 Cache-based TTA.
Inspired by the key-value storage framework introduced in TIP-Adapter [zhang2022tip], several cache-based TTA approaches [tda, dpe, boostadapter] have been proposed, which dynamically store and update high-confidence feature-label pairs during inference. The cache is organized as a priority queue, where the priority of each entry is determined by the entropy of CLIP’s predictions.
Let $\mathbf{W}_{\mathrm{CLIP}}=\left[\textbf{t}^{1},\textbf{t}^{2},\ldots,\textbf{t}^{C}\right]^{\top}$ be a stack of text embeddings for all classes. Cache item is represented as:
@@
As test samples arrive sequentially, new entries are inserted into cache slots corresponding to their pseudo-labels. When the cache exceeds its capacity, the entries with the highest entropy are evicted. During inference, a test feature $\mathbf{z}$ retrieves related cache items, and the final prediction is calculated by combining the CLIP logits with those derived from the cache:
@@
Following the DPE approach dpe, the feature for each cache slot, $\mathbf{F}_{\text{cache}}$, is computed as the mean of all stored features for that class, forming increasingly compact and representative class-specific prototypes as more data accumulates. The cache item is then constructed as:
@@
3.2 Adaptive Cache Enhancement (ACE)
Cache-based strategies demonstrate notable effectiveness in various domains. However, they are grappling with two substantial issues. First, entropy is not a reliable indicator when the data distribution changes. Second, fixed decision boundaries are not adaptive to distributional changes. To address these challenges, we introduce two specialized modules that specifically address each issue while synergistically strengthening each other, as illustrated in Figure 2. In our framework, we explore two variants of ACE to accommodate different confidence estimation strategies: ACE-Entropy and ACE-Probability. ACE-Entropy employs entropy $\mathcal{H}(p)$ as the primary metric for cache selection and threshold refinement. In contrast, ACE-Probability uses maximum class probability $\max(p_{c})$ for thresholding. Both is initialized from zero-shot statistics and updated via EMA, and both maintain computational parity, with no additional overhead.
3.2.1 Curriculum Thresholding
The task of dynamically setting class-specific thresholds in alignment with the evolving learning progress of the model is challenging. Ideally, the threshold for each class can be set specifically based on evaluation metrics, e.g., recognition accuracy or entropy as $\mathcal{T}_{t}(c)=m_{t}(c)\cdot\mathcal{T}_{0}(c)$, where $\mathcal{T}_{t}(c)$ is the adaptive threshold for class $c$ at time step $t$, and $m_{t}(c)$ denotes the corresponding metric, whether it is accuracy or entropy. This ensures that classes with poorer performance, indicated by lower accuracy rates, are designated smaller thresholds (or conversely, higher entropy leads to larger thresholds), thereby permitting a greater number of their instances to be retained in the cache during the testing phase. However, dynamically adjusting these thresholds throughout the training process requires continuous accuracy estimates at time increments $t$, which impedes training efficiency.
In this study, we introduce Curriculum Thresholding, a thresholding mechanism that evaluates the learning status of classes without introducing any additional inference steps and/or requiring additional annotations. Central to our hypothesis is the postulation that when subjected to a high threshold, the quantity of confidently classified predictions ascribed to a particular class serves as a robust indicator of the learning progress for that class. Specifically, we proceed with the following steps:
$\sigma_{t}(c)$ $=\sum_{n=1}^{N}\mathds{1}\left(\max\left(p(y|x)\right)\geq\mathcal{T}_{t-1}(c)\right)\cdot L_{p}^{c},$
or $=\sum_{n=1}^{N}\mathds{1}\left(\min\left(e(y|x)\right)\leq\mathcal{T}_{t-1}(c)\right)\cdot L_{p}^{c},$
$L_{p}^{c}$ $=\begin{cases}1,&\text{if}\ \arg\max(p(y|x))=c\\ 0,&\text{otherwise}\end{cases}$
where $\sigma_{t}(c)$ measures the number of samples confidently predicted as class $c$ at time $t$, and $p(y|x)$ and $e(y|x)$ are the probability and entropy of the model to sample $x$. We then normalize $\sigma_{t}(c)$ as $\sigma_{t}(c)\leftarrow\frac{\sigma_{t}(c)}{\max_{c}\sigma_{t}}$
@@ 3.2.1 Curriculum Thresholding
3.2.2 Online Thresholding Adaptation
@@
CLIP-ResNet-50 [radford2021learning] 58.16 21.83 51.41 56.15 33.37 44.18 40.69
@@
CoOp [zhou2022learning] 63.33 23.06 55.40 56.60 34.67 46.61 42.43
TPT [shu2022test] 60.74 26.67 54.70 59.11 35.09 47.26 43.89
DiffTPT [feng2023diverse] 60.80 31.06 55.80 58.80 37.10 48.71 45.69
TDA [karmanov2024efficient] 61.35 30.29 55.54 62.58 38.12 49.58 46.63
TPS [sui2024just] 61.47 30.48 54.96 62.87 37.14 49.38 46.36
DMN-ZS [zhang2024dual] 63.87 28.57 56.12 61.44 39.84 49.97 46.49
DPE [dpe] 63.41 30.15 56.72 63.72 40.03 50.81 47.66
CRG [zhai2025mitigatingcachenoisetesttime] 65.26 29.69 56.07 - - - -
ACE-Probability 64.41 33.50 57.50 63.95 40.15 51.75 49.37
ACE-Entropy 64.30 33.55 57.51 63.82 40.25 51.78 49.38
CLIP-ViT-B/16 [radford2021learning] 66.73 47.87 60.86 73.98 46.09 59.11 57.20
@@
CoOp [zhou2022learning] 71.51 49.71 64.20 75.21 47.99 61.72 59.28
TPT [shu2022test] 68.98 54.77 63.45 77.06 47.94 62.44 60.81
DiffTPT [feng2023diverse] 70.30 55.68 65.10 75.00 46.80 62.28 60.52
TDA [karmanov2024efficient] 69.51 60.11 64.67 80.24 50.54 65.01 63.89
TPS [sui2024just] 70.19 60.08 64.73 80.27 49.95 65.04 63.76
DMN-ZS [zhang2024dual] 72.25 58.28 65.17 78.55 53.20 65.49 63.80
DPE [dpe] 71.91 59.63 65.44 80.40 52.26 65.93 64.43
CRG [zhai2025mitigatingcachenoisetesttime] 75.01 63.67 64.66 - - - -
ACE-Probability 72.56 63.54 65.86 81.01 52.60 67.05 66.01
ACE-Entropy 72.57 63.47 65.91 81.09 52.66 67.10 66.05
@@
3.2.3 Test-Time Logit Inference and Loss Function.
We adopt DPE, the residual learning method proposed by zhang2024dual, to define textual and visual embedding prototypes as $\mathbf{t}_{c}=\frac{1}{S}\sum_{i=1}^{S}\mathcal{E}_{t}(\mathcal{P}^{c}_{i})$ and $\mathbf{v}_{c}=\frac{1}{S_{c}}\sum_{m=1}^{S_{c}}\mathbf{z}^{c}_{m}$. Here, $S$ is the number of prompts, $S_{c}\leq M$ is the total count of image features stored in the queue corresponding to class $c$, and $\mathbf{z}$ is the image features of $x$ that surpass the threshold $\mathcal{T}$. Once the two sets of multi-modal prototypes have been adjusted based on the last test sample, they are used to serve as the initial configuration for further refinement with the current test input. The preliminary outcomes derived from these steps are subsequently utilized to fine-tune the prototypes, making them adapted to $x$ as,
@@
where $\hat{\mathbf{t}}_{c}\in\mathbb{R}^{C\times d}$ and $\hat{\mathbf{v}}_{c}\in\mathbb{R}^{C\times d}$ are the learnable residual parameters that are initialized as zero and are used to optimize the prototypes for each given test input $x$. Similarly to zhang2024dual, we optimize these residual parameters to promote consistent predictions in a total of $N$ different augmented views of each given test image $x$ using the objective of unsupervised entropy minimization,
@@
where $\mathcal{A}_{n}(x)$ is an augmented view of $x$, $T_{e}$ and $T_{p}$ are the entropy threshold and probability threshold, respectively.
We adopt the alignment loss in DPE:
@@
CoOp [CoOp] 15.12 86.53 55.32 37.29 26.20 61.55 75.59 87.00 58.15 59.05 56.18
TPT [tpt] 17.58 87.02 58.46 40.84 28.33 62.69 74.88 84.49 61.46 60.82 57.66
DiffTPT [Feng2023DiverseDA] 17.60 86.89 60.71 40.72 41.04 63.53 79.21 83.40 62.72 62.67 59.85
TDA [tda] 17.61 89.70 57.78 43.74 42.11 68.74 77.75 86.18 62.53 64.18 61.03
DPE [dpe] 19.80 90.83 59.26 50.18 41.67 67.60 77.83 85.97 64.23 61.98 61.93
CRG [zhai2025mitigatingcachenoisetesttime] 18.09 90.12 57.92 51.89 46.80 71.09 75.76 85.91 63.11 63.58 62.42
@@
CoOp [CoOp] 18.47 93.70 64.51 41.92 46.39 68.71 85.30 89.14 64.15 66.55 63.88
TPT [tpt] 24.78 94.16 66.87 47.75 42.44 68.98 84.67 87.79 65.50 68.04 65.10
DiffTPT [Feng2023DiverseDA] 25.60 92.49 67.01 47.00 43.13 70.10 87.23 88.22 65.74 62.67 65.47
TDA [tda] 23.91 94.24 67.28 47.40 58.00 71.42 86.14 88.63 67.62 70.66 67.53
DPE [dpe] 28.95 94.81 67.31 54.20 55.79 75.07 86.17 91.14 70.07 70.44 69.40
CRG [zhai2025mitigatingcachenoisetesttime] 26.58 93.57 66.89 56.38 59.81 75.94 85.95 91.20 68.36 70.31 69.50